OpenStudy (anonymous):

Which of the following equations would graph a line parallel to 3y = x + 5

Nnesha (nnesha):

hint : parallel lines have same slope

OpenStudy (anonymous):

a. 3y = x + 1 b. y = x + 1 c. y = 2/3x + 1 d. 3/2y = x + 1 I chose a

Nnesha (nnesha):

yes that is right because then you have same slope which is 1/3

OpenStudy (anonymous):

I knew I was right XD

Nnesha (nnesha):

yes :) :)